OpenStudy (anonymous):

In a certain clothing store, 6 shirts and 3 ties cost $79.50, and 3 shirts and 2 ties cost $41. Determine the cost of each tie.

OpenStudy (anonymous):

can i have the answer im in summer school is it a)$2.50 b)$3.50 c)$4.50 which one please

OpenStudy (anonymous):

6+3=$79.50 + 3+2=$41

OpenStudy (anonymous):

6s+3t=$79.50 3s+2t=$41

OpenStudy (anonymous):